# Security of Language Models for Code: A Systematic Literature Review

### v1

@@ 7. Answer to RQ4: Experimental Setting and Evaluation
7.4. RQ3.4: Artifact Accessibility
7.4.1. Open-Source Code Repository in CodeLMs Studies
@@
Table 12. Replication open-source code repository links provided in CodeLM studies
Year Technique URL
2020 DAMP (Yefet et al., 2020) https://github.com/tech-srl/adversarial-examples
2020 Bielik et al. (Bielik and Vechev, 2020) https://github.com/eth-sri/robust-code
2020 MHM (Zhang et al., 2020a) https://github.com/Metropolis-Hastings-Modifier/MHM
2021 Severi et al.(Severi et al., 2021) https://github.com/ClonedOne/MalwareBackdoors
2021 Srikant et al.(Srikant et al., 2021) https://github.com/ALFA-group/adversarial-code-generation
2021 Pour et al.(Pour et al., 2021) https://github.com/MaryamVP/Guided-Mutation-ICST-2021
2021 Ferretti et al.(Ferretti and Saletta, 2021) https://github.com/Martisal/adversarialGE
2021 Anand et al.(Anand et al., 2021b) https://github.com/mtensor/neural_sketch
2021 ACCENT (Zhou et al., 2022) https://github.com/zhangxq-1/ACCENT-repository
2022 Ramakrishnan et al.(Ramakrishnan and Albarghouthi, 2022) https://github.com/tech-srl/code2seq
2022 Wan et al.(Wan et al., 2022) https://github.com/CGCL-codes/naturalcc
2022 CARROT (Zhang et al., 2022a) https://github.com/SEKE-Adversary/CARROT
2022 ALERT (Yang et al., 2022) https://github.com/soarsmu/attack-pretrain-models-of-code
2022 RoPGen (Li et al., 2022a) https://github.com/RoPGen/RoPGen
2022 Ramakrishnan et al.(Ramakrishnan and Albarghouthi, 2022) https://github.com/goutham7r/backdoors-for-code
2022 Henkel et al.(Henkel et al., 2022) https://github.com/jjhenkel/averloc
2022 SPACE (Li et al., 2022b) https://github.com/EricLee8/SPACE
2023 Cotroneo et al.(Cotroneo et al., 2024) https://github.com/dessertlab/Targeted-Data-Poisoning-Attacks
2023 BADCODE (Sun et al., 2023a) https://github.com/wssun/BADCODE
2023 Li et al.(Li et al., 2023b) https://github.com/Lyz1213/Backdoored_PPLM
2023 PELICAN (Zhang et al., 2023d) https://github.com/ZhangZhuoSJTU/Pelican
2023 CodeAttack (Jha and Reddy, 2023) https://github.com/reddy-lab-code-research/CodeAttack
2023 Zhang et al.(Zhang et al., 2023b) https://github.com/18682922316/RNNS-for-code-attack
2023 CODA (Tian et al., 2023) https://github.com/tianzhaotju/CODA
2023 RoPGen (Li et al., 2022a) https://github.com/RoPGen/RoPGen
2023 SVEN (He and Vechev, 2023) https://github.com/eth-sri/sven
2024 AFRAIDOOR (Yang et al., 2024c) https://github.com/yangzhou6666/adversarial-backdoor-for-code-models
2024 TROJANPUZZLE (Aghakhani et al., 2024) https://github.com/microsoft/CodeGenerationPoisoning
2024 CodePoisoner(Li et al., 2024b) https://github.com/LJ2lijia/CodeDetector
2024 DeCE(Yang et al., 2024d) https://github.com/NTDXYG/DeCE
2024 RADAR(Yang et al., 2024e) https://github.com/NTDXYG/RADAR
@@ 8.1. Challenges and Opportunities in Attacks against CodeLMs
Stealthiness of Backdoor Triggers. In backdoor attacks on CodeLMs, the stealthiness of triggers is a key focus for attackers. They continually explore more covert trigger designs, evolving from early methods like dead code (Ramakrishnan and Albarghouthi, 2022; Wan et al., 2022) to variable/function names (Sun et al., 2023a; Li et al., 2024b), and even adaptive triggers (Yang et al., 2024c), aiming to inject increasingly stealthy triggers into code. However, comprehensively evaluating the stealthiness of triggers remains a challenging task. Current research methods typically focus on specific aspects, such as syntactic or semantic visibility, or rely on human experiments. For instance, BadCode (Sun et al., 2023a) outperforms Wan et al. (Wan et al., 2022) in human experiment evaluations, while A FRAIDOOR (Yang et al., 2024c) uses automated evaluations to demonstrate that adaptive triggers are more stealthy. Nonetheless, these methods do not cover all possible detection dimensions, leaving room for improvement in both evaluation metrics and techniques.
@@ Acknowledgments
This work is partially supported by the National Natural Science Foundation of China under grant numbers 61932012 and 62372228. Weisong Sun, Zhenpeng Chen, and Yang Liu are supported by the National Research Foundation, Singapore, and DSO National Laboratories under the AI Singapore Programme (AISG Award No: AISG2-GC-2023-008), as well as the National Research Foundation, Singapore, and the Cyber Security Agency of Singapore under the National Cybersecurity R&D Programme (NCRP25-P04-TAICeN). Any opinions, findings, conclusions, or recommendations expressed in this paper are those of the author(s) and do not reflect the views of the National Research Foundation, Singapore, or the Cyber Security Agency of Singapore.

### v2

@@ 7. Answer to RQ4: Experimental Setting and Evaluation
7.4. RQ4.4: Artifact Accessibility
Table 13. Replication open-source code repository links provided in CodeLM studies.check
Technique Year Type URL
Ramakrishnan et al. (Ramakrishnan and Albarghouthi, 2022) 2022 Data poisoning Pre-training defense https://github.com/goutham7r/backdoors-for-code
Schuster et al. (Schuster et al., 2021) 2021 Data poisoning Model poisoning https://github.com/ClonedOne/MalwareBackdoors
Severi et al.(Severi et al., 2021) 2021 Data poisoning https://github.com/ClonedOne/MalwareBackdoors
Wan et al. (Wan et al., 2022) 2022 Data poisoning https://github.com/CGCL-codes/naturalcc
BadCode (Sun et al., 2023) 2023 Data poisoning https://github.com/wssun/BADCODE
TrojanSQL (Zhang et al., 2023e) 2023 Data poisoning https://github.com/jc-ryan/trojan-sql
Cotroneo et al. (Cotroneo et al., 2024) 2023 Data poisoning https://github.com/dessertlab/Targeted-Data-Poisoning-Attacks
Li et al. (Li et al., 2024b) 2024 Data poisoning Pre-training defense https://github.com/LJ2lijia/CodeDetector
AFRAIDOOR (Yang et al., 2024c) 2024 Data poisoning https://github.com/yangzhou6666/adversarial-backdoor-for-code-models
TrojanPuzzle 2024 Data poisoning https://github.com/microsoft/CodeGenerationPoisoning
FDI (Sun et al., 2024a) 2024 Data poisoning https://github.com/v587su/FDI
LateBA (Yi et al., 2024) 2024 Data poisoning https://github.com/yxy-whu/LateBA
Li et al. (Li et al., 2023c) 2023 Model poisoning https://github.com/Lyz1213/Backdoored_PPLM
PELICAN (Zhang et al., 2023b) 2023 Model poisoning https://github.com/ZhangZhuoSJTU/Pelican
DeCE (Yang et al., 2024f) 2024 In-training defense https://github.com/NTDXYG/DeCE
DAMP (Yefet et al., 2020) 2020 White-box attack https://github.com/tech-srl/adversarial-examples
Bielik et al. (Bielik and Vechev, 2020) 2020 White-box attack https://github.com/eth-sri/robust-code
Srikant et al. (Srikant et al., 2021) 2021 White-box attack https://github.com/ALFA-group/adversarial-code-generation
Quiring et al. (Quiring et al., 2019) 2019 Black-box attack https://github.com/EQuiw/code-imitator
MHM (Zhang et al., 2020a) 2020 Black-box attack https://github.com/Metropolis-Hastings-Modifier/MHM
Pour et al. (Pour et al., 2021) 2021 Black-box attack https://github.com/MaryamVP/Guided-Mutation-ICST-2021
Ferretti et al. (Ferretti and Saletta, 2021) 2021 Black-box attack https://github.com/Martisal/adversarialGE
Anand et al. (Anand et al., 2021b) 2021 Black-box attack https://github.com/mtensor/neural_sketch
ACCENT (Zhou et al., 2022) 2021 Black-box attack https://github.com/zhangxq-1/ACCENT-repository
CARROT (Zhang et al., 2022a) 2022 Black-box attack Input modification https://github.com/SEKE-Adversary/CARROT
ALERT (Yang et al., 2022) 2022 Black-box attack https://github.com/soarsmu/attack-pretrain-models-of-code
Li et al. (Li et al., 2022a) 2022 Black-box attack Input modification https://github.com/RoPGen/RoPGen
CodeAttack (Jha and Reddy, 2023) 2023 Black-box attack https://github.com/reddy-lab-code-research/CodeAttack
RNNS (Zhang et al., 2023a) 2023 Black-box attack https://github.com/18682922316/RNNS-for-code-attack
CODA (Tian et al., 2023) 2023 Black-box attack https://github.com/tianzhaotju/CODA
CodeBERT-Attack (Zhang et al., 2024c) 2023 Black-box attack https://gitfront.io/r/DrLC5417/5pBsYXgKinB3/CodeBERT-Attack/tree/codebert/
RADAR (Yang et al., 2024g) 2024 Black-box attack Additional Models https://github.com/NTDXYG/RADAR
STRUCK (Zhang et al., 2023d) 2024 Black-box attack https://github.com/zhanghaha1707/STRUCK
CodeTAE (Yang et al., 2024a) 2024 Black-box attack https://github.com/yyl-github-1896/CodeTAE
MOAA (Zhou et al., 2024) 2024 Black-box attack https://github.com/COLA-Laboratory/MOAA
MalwareTotal (He et al., 2024) 2024 Black-box attack https://github.com/Optimus-He/MalwareTotal
Henkel et al. (Henkel et al., 2022) 2022 Input modification https://github.com/jjhenkel/averloc
Cristina et al. (Improta et al., 2023) 2023 Input modification https://github.com/dessertlab/Robustness-of-AI-Offensive-Code-Generators
SVEN (He and Vechev, 2023) 2023 Model modification https://github.com/eth-sri/sven
SPACE (Li et al., 2022b) 2022 Additional models https://github.com/EricLee8/SPACE
@@ 8.1. Challenges and Opportunities in Attacks against CodeLMs
Stealthiness of Backdoor Triggers. In backdoor attacks on CodeLMs, the stealthiness of triggers is a key focus for attackers. They continually explore more covert trigger designs, evolving from early methods like dead code (Ramakrishnan and Albarghouthi, 2022; Wan et al., 2022) to variable/function names (Sun et al., 2023; Li et al., 2024b), and even adaptive triggers (Yang et al., 2024c), aiming to inject increasingly stealthy triggers into code. However, comprehensively evaluating the stealthiness of triggers remains a challenging task. Current research methods typically focus on specific aspects, such as syntactic or semantic visibility, or rely on human experiments. For instance, BadCode (Sun et al., 2023) outperforms Wan et al. (Wan et al., 2022) in human experiment evaluations, while AFRAIDOOR (Yang et al., 2024c) uses automated evaluations to demonstrate that adaptive triggers are more stealthy. Nonetheless, these methods do not cover all possible detection dimensions, leaving room for improvement in both evaluation metrics and techniques.
@@
Exploring the Security of CodeLLMs from Other Attack Perspectives. In addition to backdoor and adversarial attack threats, recent research has also explored other types of attack threats against CodeLLMs, such as privacy leakage (Niu et al., 2023; Yang et al., 2024e; Al-Kaswan et al., 2024; Huang et al., 2023), membership inference attacks (Yang et al., 2024d; Al-Kaswan et al., 2024; Zhang et al., 2024b) and watermark attack (Xiao et al., 2025). Exploring new attack threats is crucial for a deeper understanding of the security of CodeLLMs. By studying these attack methods, researchers can identify potential vulnerabilities that the model may face in real-world applications, reveal its weaknesses in various scenarios, and better understand the model’s behavior. Therefore, conducting more research on different types of attacks will help both academia and industry more comprehensively assess the security of CodeLLMs and provide a foundation for future security research.
@@ 8.2. Challenges and Opportunities in Defenses against Attacks on CodeLMs
Defending Against Emerging Security Threats in CodeLLMs. With the emergence of new security threats such as privacy leakage and membership inference attacks, it has become crucial to extend existing defense techniques to address these risks. Although previous studies have proposed defense techniques to enhance the privacy protection of CodeLLMs (Fried et al., 2023; Allal et al., 2023; Li et al., 2023a), these defenses still face challenges in the face of the continuously emerging attack threats. Therefore, developing defense strategies that effectively address the evolving threats is essential to maintaining the security of CodeLLMs in real-world applications.
@@
9. Threats to Validity
Potential Impact of Database Selection on the Inclusion of Relevant Papers. We conduct automated searches across six widely used databases, including Google Scholar, ACM Digital Library, IEEE Xplore, Springer, DBLP, and ArXiv, all of which are standard and commonly used online engines or databases. One potential issue is that relevant papers may be omitted if the search relies solely on these databases. To mitigate this impact, we also employ both forward and backward snowballing search methods to further enhance the search results. Through snowballing, we identify papers with transitive dependencies and expand our paper collection. Specifically, backward snowballing helps us uncover relevant research in the reference lists of the collected papers, while forward snowballing tracks subsequent studies that cite the papers we have collected, further extending the scope of our search. This approach ensures comprehensive coverage of relevant literature, reducing the risk of omissions.
Potential Impact of Non-peer-reviewed Papers on the Reliability of the Conclusions. We include ArXiv in our search databases, considering that some researchers are inclined to disclose the latest techniques on ArXiv in advance. As a result, the final selection of papers may include non-peer-reviewed papers. One potential impact is that these non-peer-reviewed papers may affect the reliability of our research conclusions. To mitigate this impact, we ensure the quality of the selected papers. We apply five exclusion criteria to rigorously filter all the collected papers. For the final set of selected papers, the first two authors review them thoroughly and multiple times to ensure that each paper has a clear research motivation, provides detailed descriptions of the techniques, presents comprehensive experimental setups, and clearly confirms the experimental findings, thus ensuring that low-quality papers are excluded from our study. Additionally, our research conclusions primarily encompass the development trends in the field (Section 3) and summaries of various research directions (Sections 4 to 7). These conclusions do not include specific technique comparisons between the papers. Therefore, the inclusion of non-peer-reviewed papers does not significantly impact the reliability of our conclusion.
Potential Bias in Study Selection. In the study selection process, we establish five inclusion and exclusion criteria to filter the papers, combining automated and manual procedures. In the first and second phases of the study selection, we use automated filtering methods, as the inclusion and exclusion criteria in these stages are easily aligned with relevant fields in the BibTeX records. However, in the third to fifth phases, due to the difficulty in directly correlating the inclusion and exclusion criteria with specific fields in the BibTeX records, the risk of mislabeling papers is higher. Therefore, we use manual filtering. During this process, we carefully review the papers multiple times to mitigate potential biases introduced by manual filtering. However, manual filtering may still be influenced by the researchers’ subjective judgment biases, which could affect the accuracy of the quality assessment of papers. To address these concerns, the first two authors carefully review the selected papers multiple times, focusing on the abstract, introduction, related work, experiments, and conclusion sections, to fully understand the research objectives and categorization of each paper. This step aims to improve the accuracy of paper selection, minimize omissions, and reduce the potential bias in the study selection process.
@@ Acknowledgments
This research is supported by the National Key Research and Development Program of China (2024YFF0908001) and the National Natural Science Foundation of China (U24A20337 and 62372228), the National Research Foundation, Singapore, and DSO National Laboratories under the AI Singapore Programme (AISG Award No: AISG2-GC-2023-008), as well as the National Research Foundation, Singapore, and the Cyber Security Agency of Singapore under the National Cybersecurity R&D Programme (NCRP25-P04-TAICeN). Any opinions, findings, conclusions, or recommendations expressed in this paper are those of the author(s) and do not reflect the views of the National Research Foundation, Singapore, or the Cyber Security Agency of Singapore.